A rose garden is planted on a square plot of land. A wooden fence around the garden’s perimeter has a total length of 320 feet. What is the approximate length of the diagonal of the garden, to the nearest foot?
Answer by satyareddy22(84)

Let a be the side of a square plot
perimeter = 320 feet
4a = 320
a=320/4 = 80
a=80
by Pythagorean theorem
a^2 + a^2 =(diagonal)^2
80^2 + 80^2 = (diagonal)^2
6400 + 6400 =(diagonal)^2
12800 = (diagonal)^2
diagonal = sq rt(12800) = 11.314
diagonal = 11.314